OpenStudy (anonymous):

Solve the following system of equations: −2x + y = 1 −4x + y = −1 Answers (multiple choice) 1. (3, 1) 2. (−1, 3) 3. (−1, −3) 4. (1, 3)

OpenStudy (anonymous):

answer is (1,3) x being 1 and y being 3

OpenStudy (anonymous):

since you have it as a multiple choice I just pluged the numbers in and the only one working was (1,3) -2(1)+(3)=1 -4(1)+(3)=-1

OpenStudy (anonymous):

THX

OpenStudy (anonymous):

anytime :)